Let X be a discrete random variable such that X = 0 with probability 0.5 and X = 1 with probability 0.5. Let Y be a discrete random variable such that Y = 0 when X = 1 and Y = 1 when X = 0. What is the mean and variance of X + Y?


1
Expert's answer
2022-01-26T09:50:54-0500

If "X=0," then "X+Y=0+1=1."

If "X=1," then "X+Y=1+0=1."

Then "P(X+Y=1)=1."


"mean=E(X+Y)=1"

"Var(X+Y)=E((X+Y)^2)-(E(X+Y))^2=0"